OpenStudy (anonymous):

Solve the quadratic equation x2 + 16x = -64 by factoring. A.{2, 4} B.8 C.{-8, 8} D.-8

OpenStudy (superdavesuper):

Try adding 64 to both sides of the equation and you will get...

OpenStudy (anonymous):

x2 + 16x = -64 +64 +64 x2nd+16x+64=0 (x+8)(x+8)=0 (x+8)=0 (x+8)=0 x=-8 x=-8

OpenStudy (anonymous):

make sure to set the problem = to 0 also - or + any like terms then find what times what would give you 64 but add up to 16. you also want to make sure that if its + and + such as +16x+64 then the number multiplied would be a + number but it if its - and + then one number would be - and the other would be +.

OpenStudy (anonymous):

after you find what the number or numbers are, set them equal to 0 like i did there (x+8)=0 (x+8)=0 the answer would be the opp of what there sign would be so x=-8 and x=-x

OpenStudy (anonymous):

oops i mean x=-8 and x=-8